OpenStudy (anonymous):

Find the rate per unite of time. 60 miles in 30 minutes, $100 in 8 hours

OpenStudy (anonymous):

60/30=2/1 or 2 100/8=25/2

OpenStudy (anonymous):

rate is simply a ratio

OpenStudy (anonymous):

o.k and $100 in 8hours?]